The percent distribution of live multiple-delivery
Answer #1

a) From the pie chart we see that there are 37.2% mothers in between age 30 - 34 and 24.6% in between 35 - 39

Hence the required probability is 0.372+0.246=0.618

That is 61.8%

b) Since P(not A) =1-P(A)

So the required probability is 1-0.618=0.382

That is 38.2%

c) Since there are only 3.7% mothers of age greater than 45,

So the required probability is 1-0.037=0.963

That is 96.3%

d) As there are 1.3% mothers of age less than 20

So the required probability is 1-0.013=0.987

That is 98.7%